# HD 110067 c has an aligned orbit

### v1

@@ 4 Analysis
The obtained stellar radial velocities clearly show the anomaly produced by the R-M effect (Fig. 2). To measure the projected stellar obliquity of the system ($\lambda$), we fit the radial velocities with a composite model, which includes a Keplerian orbital component, as well as the R-M anomaly. This model is implemented in the ARoMEpy [^3] https://github.com/esedagha/ARoMEpy (Sedaghati et al., 2023) package, which utilises the Radvel python module (Fulton et al., 2018) for the formulation of the Keplerian orbit. ARoMEpy is a Python implementation of the R-M anomaly described in the ARoME code (Boué et al., 2013). We have used the R-M effect function defined for radial velocities determined through the cross-correlation technique in our code. We have set wide uniform priors for the radial velocity semi-amplitude $K$ and the systemic velocity $\gamma$. In the R-M effect model, we fix the following parameters to values reported in the literature: orbital period $P$ and the eccentricity $e$ to a circular orbit. The parameter $\sigma$, which is the width of the Cross-correlation function (CCF), is measured on the data and fixed. Furthermore, we use the ExoCTK [^4] https://github.com/ExoCTK/exoctk tool to compute the quadratic limb-darkening coefficients with ATLAS9 model atmospheres (Castelli & Kurucz, 2003) in the wavelength range of the ESPRESSO instrument (380-788 nm). We set Gaussian priors from the discovery paper (Luque et al., 2023) on the following parameters during the fitting procedure: central transit time $T_{C}$, orbital inclination $i$, the scaled semi-major axis $a/R_{\rm{s}}$ and the planet-to-star radius ratio $R_{\rm{p}}/R_{\rm{s}}$. Uniform priors were set on the projected stellar rotational velocity $\nu\,\sin i_{*}$ and sky-projected angle between stellar rotation axis and normal of the orbital plane $\lambda$. To obtain the best fitting values of the parameters we employed three independent Markov Chain Monte Carlo (MCMC) simulations each with 250 000 steps, burning the first 50 000. The results of the MCMC simulations are presented in Table 3 and in Figs. 2 and 4.
@@ 5 An aligned system with small TTVs
From our analysis, we find a well aligned orbit for HD 110067 c with sky projected obliquity $\lambda=6^{+24}_{-26}$ degrees. Such a high state of alignment is indicative of the fact that the architecture of the system has been reached through convergent migration without any major disruptive events (making the reasonable assumption that all the planets share the same value of $\lambda$ – see Fig. 6). Given the tidal circularisation timescale of planet c ($\tau_{\rm{circ}}\gg$ 14 Gyr, Goldreich & Soter, 1966), we can also rule out any events exciting the orbit of planet c that would later be realigned, thereby confirming the quiescent evolution of the system.
@@
Figure 3: Estimated transit time variations (TTV) for all the planets in the system, assuming nominal masses from Luque et al. (2023). The dashed vertical line marks the observing time with ESPRESSO.
@@
Constraining the amplitude of the TTVs is rather difficult with such a weak constraint on the planetary masses. Nevertheless, we have run numerical simulations using REBOUND (Rein & Liu, 2012, see Figs. 3 and 7) and indeed, since the published ephemeris, TTVs up to tens of hours could be expected. In this sense, having caught the transit with such a small TTV appears quite surprising, even if this is consistent with the reported nondetection of TTVs in two years of TESS observations (Luque et al., 2023). The masses in the system are not well known, however – for some of them, only upper limits exist. This will lead to a rather large range of variations in the expected TTVs for planet c as shown in Fig. 8. Some combination of masses would explain the small amount of TTVs since the previous ephemeris. However, as the solution is clearly degenerate, more transit times are definitively needed to confirm this. Such TTVs need to be taken into account when planning future atmospheric follow-up, during, e.g., JWST Cycle 4 or Ariel to avoid wasting valuable telescope time.
@@ Appendix B Additional figures
Figure 6: Result of a REBOUND simulation in which, initially, the most external planet was inclined by 10 degrees with respect to the others. The left panel shows a view from above, while the right panel illustrates the edge-on view. The plots show the result after 5 000 years, clearly illustrating that all planets are now inclined. Moreover, the excursion of the planets around their mean semi-major axis is now higher, and this would thus lead to much larger TTVs.
Figure 7: Orbits of the planets in the system, integrated over five years and showing how they oscillate around a mean value. Such variations are the cause of the transit-timing variations. The observer is on the top of the figure.
Figure 8: Estimated transit time variations for planet c in the system, for various masses of the planets, as indicated in the legend. The dashed vertical line marks the observing time with ESPRESSO.

### v2

@@ 4 Analysis
The obtained stellar radial velocities clearly show the anomaly produced by the R-M effect (Fig. 2). To measure the projected stellar obliquity of the system ($\lambda$), we fit the radial velocities with a composite model, which includes a Keplerian orbital component, as well as the R-M anomaly. This model is implemented in the ARoMEpy [^3] https://github.com/esedagha/ARoMEpy (Sedaghati et al., 2023) package, which utilises the Radvel python module (Fulton et al., 2018) for the formulation of the Keplerian orbit. ARoMEpy is a Python implementation of the R-M anomaly described in the ARoME code (Boué et al., 2013). We have used the R-M effect function defined for radial velocities determined through the cross-correlation technique in our code. We have set wide uniform priors for the radial velocity semi-amplitude $K$ and the systemic velocity $\gamma$. In the R-M effect model, we fix the following parameters to values reported in the literature: orbital period $P$ and the eccentricity $e$ to a circular orbit. The parameter $\sigma$, which is the width of the Cross-correlation function (CCF) and represents the effects of the instrumental and turbulent broadening, is measured on the data and fixed to $\sigma=6.60$ km/s. Furthermore, we use the ExoCTK [^4] https://github.com/ExoCTK/exoctk tool to compute the quadratic limb-darkening coefficients with ATLAS9 model atmospheres (Castelli & Kurucz, 2003) in the wavelength range of the ESPRESSO instrument (380-788 nm). The used limb-darkening coefficients are $u_{1}=0.564$ and $u_{2}=0.145$. We set Gaussian priors from the discovery paper (Luque et al., 2023) on the following parameters during the fitting procedure: central transit time $T_{C}$, orbital inclination $i$, the scaled semi-major axis $a/R_{\rm{s}}$ and the planet-to-star radius ratio $R_{\rm{p}}/R_{\rm{s}}$. Uniform priors were set on the projected stellar rotational velocity $\nu\,\sin i_{*}$ and sky-projected angle between stellar rotation axis and normal of the orbital plane $\lambda$. To obtain the best fitting values of the parameters we employed three independent Markov Chain Monte Carlo (MCMC) simulations each with 250 000 steps, burning the first 50 000. The results of the MCMC simulations are presented in Table 3 and in Figs. 2 and 4.
@@ 5 An aligned system with small TTVs
From our analysis, we find a well aligned orbit for HD 110067 c with sky projected obliquity $\lambda=6^{+24}_{-26}$ degrees. Such a high state of alignment is indicative of the fact that the architecture of the system has been reached through convergent migration without any major disruptive events (making the reasonable assumption that all the planets share the same value of $\lambda$). We ran a numerical simulation with one planet inclined and such system does not lead to a set of six transiting exoplanets. Given the tidal circularisation timescale of planet c ($\tau_{\rm{circ}}\gg$ 14 Gyr, Goldreich & Soter, 1966), we can also rule out any events exciting the orbit of planet c that would later be realigned, thereby confirming the quiescent evolution of the system.
@@
Figure 3: Estimated transit time variations for planet c in the system, for various masses of the planets, as indicated in the legend. The dashed vertical line marks the observing time with ESPRESSO. The dotted horizontal lines mark no TTV (black) and a 19 min TTV (red), respectively.
@@
Constraining the amplitude of the TTVs is rather difficult with such a weak constraint on the planetary masses (Lammers & Winn, 2024). Nevertheless, we have run numerical simulations using REBOUND (Rein & Liu, 2012) and indeed, since the published ephemeris, TTVs up to tens of hours could be expected. In this sense, having caught the transit with such a small TTV appears quite surprising, even if this is consistent with the reported nondetection of TTVs in two years of TESS observations (Luque et al., 2023). The masses in the system are not well known, however – for some of them, only upper limits exist. This will lead to a rather large range of variations in the expected TTVs for planet c as shown in Fig. 3. Some combination of masses would explain the small amount of TTVs since the previous ephemeris. However, as the solution is clearly degenerate, more transit times are definitively needed to confirm this. Such TTVs need to be taken into account when planning future atmospheric follow-up, during, e.g., JWST Cycle 4 or Ariel to avoid wasting valuable telescope time.